OpenStudy (anonymous):

-7z - 7 is less than or the same as - 5z + 12

OpenStudy (anonymous):

plug in z =0 see what u get

OpenStudy (anonymous):

it says solve the salloing equations write the variables above their valures to find the name of a famous volcano a+18= -32??

OpenStudy (anonymous):

Nope it was wrong.. and whatare you talking about madhatter??

OpenStudy (anonymous):

i need help! sheg???? it says solve the salloing equations write the variables above their valures to find the name of a famous volcano a+18= -32??

OpenStudy (anonymous):

yes music wasted -7z-7 is less than -5z+12

OpenStudy (anonymous):

sheg ur a real help :) please help me

OpenStudy (anonymous):

ok madhatter let me see ur problem

OpenStudy (anonymous):

it says solve the salloing equations write the variables above their valures to find the name of a famous volcano a+18= -32??

OpenStudy (anonymous):

Okay but your supposed to solve for z... what is z?